OpenStudy (anonymous):

Determine whether the given points are on the graph of the equation y = 8x − 6; a)(0, 6), b)(1/8,5) c)(1,2)

OpenStudy (anonymous):

a 6=(8*0)-6 doesnt work point not on line b 5=(1/8*8)-6 works so point is on line c 2=8(1)-6 works so point is on the line

OpenStudy (anonymous):

oops actually b doesn't work 1-6 isn't 5 :P

OpenStudy (rosey):

u should b a fan of me :)